# Isocurvature Induced Gravitational Waves at Pulsar Timing Arrays

### v1

@@ 1 Introduction
Gravitational waves (GWs) serve as another powerful probe of early-Universe physics. Because they free-stream, GWs can access information from much earlier times than the CMB photons. This prospect drives the continued, worldwide development of GW observatories designed to measure a broad range of frequencies. In particular, Pulsar Timing Arrays (PTA) provide a unique probe of GWs in the nano-Hertz frequency range. Several experiments such as NANOGrav [NANOGrav:2023gor], EPTA/InPTA [EPTA:2023fyk], PPTA [Reardon:2023gzh], CPTA [Xu:2023wog] have recently observed a stochastic GW background and derived constraints for new physics. One possible source of GW that might contribute to PTA observations is called scalar induced GWs (SIGW) [^1] See Ref. [Domenech:SIGWRe] for a review. [Matarrese:1992rp, Matarrese:1993zf, Matarrese:1997ay, Ananda:2006af, Baumann:2007zm]: primordial scalar perturbations can source GWs at the second order in the perturbation theory. This framework has been applied to adiabatic perturbations and a certain type of isocurvature in cold dark matter (CDM) [Domenech:2021and, Domenech:2023jve, Chen:2024twp, Yuan:2024qfz, He:2024luf, Luo:2025lgr, Han:2025wlo, Ali:2025xvj, Zeng:2025tno, Domenech:2025ffb, Papanikolaou:2025ddc, Garcia:2025yit, Yuan:2025seu]. In this work, we present a more comprehensive study of isocurvature induced GWs including four types of isocurvature: cold dark matter density isocurvature (CDI), baryon density isocurvature (BDI), neutrino density isocurvature (NDI) and free-streaming dark radiation density isocurvature (DRDI). We emphasize that, for the frequency range relevant for PTA observations, the modes entered the horizon before SM neutrino decoupling (when the temperature of the Universe is around MeV), and thus neutrinos are tightly coupled to the SM bath. Therefore, the initial conditions for NDI are qualitatively different from the standard ones where neutrinos are treated as free-streaming radiation. In this work, we derive novel initial conditions for all isocurvature modes with coupled neutrinos, and calculate the evolution of density and metric perturbations numerically. We find that the GW spectrum from NDI with coupled neutrinos is the same as that of CDI or BDI up to an overall coefficient, which leads to an interesting ‘conversion’ of isocurvature between radiation and matter sectors. Furthermore, DRDI results are dramatically different from those in the previously studied CDI case because of the presence of anisotropic stress. The size of DRDI induced GWs is significantly enhanced for high $k$ modes compared to that of CDI because the ratio of dark radiation (DR) density to the total bath is not suppressed at early times unlike the matter density. Therefore, PTA opens a new window to test BSM physics and places novel constraints via isocurvature induced GWs.
@@
To derive general constraints that can be applied to a broad class of models, we use two parameterizations of the primordial isocurvature power spectrum: a delta function and a broken power law [Buckley:2025zgh]. Using the general formula for the source term of induced GWs, we obtain the spectrum of GWs and apply the data from NANOGrav collaboration to place constraints on the isocurvature power spectrum. Our results set stringent limits for wavenumbers around $10^{6}\,\text{Mpc}^{-1}$, which are complementary to other cosmological constraints on large scales.
@@ 2 Isocurvature Initial Conditions Before Neutrino Decoupling
To study perturbations in cosmology [Kodama:1986fg, Kodama:1986ud, Ma:1995ey], we start with the perturbed FRW metric in the conformal Newtonian gauge,
@@
Now we move on to study different types of perturbation in cosmology. The gauge invariant curvature perturbation can be written in the Newtonian gauge as [Wands:2000dp]
@@
where $R_{i}\equiv\bar{\rho}_{i}/\bar{\rho}_{r}$ with $\bar{\rho}_{r}$ being the total background radiation density and $i$ denotes species in the radiation bath. The gauge invariant isocurvature perturbation, or the relative entropy perturbation is defined as
${\cal S}_{\alpha\beta}=3(\zeta_{\alpha}-\zeta_{\beta})~.$ (2.8)
@@
All other solutions are called isocurvature initial conditions, which have vanishing curvature perturbation $\zeta=0$ but at least one non-zero $\cal S_{\alpha\beta}$. Since there are multiple species in the Universe, we can define different isocurvature initial conditions by different choices of non-vanishing $\cal S_{\alpha\beta}$. In this paper, we choose the adiabatic SM radiation bath $\gamma_{*}$ as the reference and we consider only one non-vanishing $\mathcal{S}_{\alpha\gamma_{*}}$ in each case as $\tau\to 0$. [^2] Since $\mathcal{S}_{\alpha\gamma_{*}}$ is always the same as $\mathcal{S}_{\alpha\gamma}$, it is equivalent to choose the photon bath as the reference. For example, NDI is defined by setting $\mathcal{S}_{\alpha\gamma_{*}}=0$ except for $\mathcal{S}_{\nu\gamma_{*}}\neq 0$. Similarly, we can define BDI ($\mathcal{S}_{b\gamma_{*}}\neq 0$), CDI ($\mathcal{S}_{c\gamma_{*}}\neq 0$), and DRDI ($\mathcal{S}_{\textrm{dr}\gamma_{*}}\neq 0$). We note that the definition of isocurvature conditions is not unique, which depends on the choice of non-vanishing $\cal S_{\alpha\beta}$. Another commonly studied choice is based on $\mathcal{S}_{\alpha r}$ which involves the perturbation of the radiation bath [Bucher:1999re]. We find this choice matches our definition if the isocurvature component is not in the radiation bath like BDI and CDI, but differs for NDI and DRDI. Since $\zeta_{r}$ in the NDI case contains neutrino perturbations that have isocurvature, setting $\mathcal{S}_{br/cr}=0$ does not respect the adiabaticity of baryons and CDM. Therefore, for the study of new physics that generate isocurvature in certain species and the rest remain adiabatic, it is more convenient to use the definition based on $\mathcal{S}_{\alpha\gamma_{*}}$ instead of $\mathcal{S}_{\alpha r}$.
@@
As mentioned before, we separate $\nu$ from $\gamma_{*}$ for NDI. Setting $\mathcal{S}_{\nu\gamma_{*}}=3/4$, [^3] This is the convention commonly used for radiation isocurvature, which ensures the leading isocurvature perturbation (e.g., $\delta_{\nu}$ in NDI) is normalized close to unity [Bucher:1999re]. we get perturbations in NDI as
@@
$\theta_{c}$ $=$ $\frac{1}{3}\theta_{b}=\frac{1}{3}\theta_{\gamma_{*}}=-\frac{1}{24}R_{\nu}\omega k^{2}\tau^{2}.$ (2.13)
@@
$\theta_{c}$ $=$ $-\frac{(1-R_{\rm dr})R_{\rm dr}}{15+4R_{\rm dr}}k^{2}\tau~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~(\textrm{DRDI})$
@@
We show the time evolution of $\Psi$ and $\Phi$ for different cases in the left panel of figure 1. For illustrative purposes, we just show the AD, CDI and DRDI cases. Since initial conditions for CDI is qualitatively similar to BDI and NDI up to an overall rescaling as mentioned before, these choices covers all relevant features. As seen from figure 1, $\Psi=\Phi\propto C$ for AD and $\Psi=\Phi\propto\omega\tau$ for CDI at superhorizon. They both become damped oscillations after entering the horizon. Around horizon crossing ($k\tau\sim 1$), the CDI case is suppressed by $\omega/k$ (or $k_{\rm eq}/k$ because $k_{\rm eq}=\sqrt{2}\omega$) for large $k$. Therefore we rescale the CDI case by the inverse of this factor. The DRDI case, however, has qualitatively different $\Phi$ and $\Psi$ evolutions due to the presence of anisotropic stress from free-streaming DR. In DRDI, $\Phi,\Psi,\Phi-\Psi\propto R_{\rm dr}$ (see Eq. (2.16)) at superhorizon. They both start damped oscillations after horizon entry and become the same deep inside the horizon due to the anisotropic stress being suppressed (see Eq. (2.15)). Moreover, we present the evolution of the curvature perturbation $\zeta$ in the right panel of figure 1. $\zeta$ of AD starts at one by definition and damps after horizon entry. For isocurvature cases, $\zeta$ starts at zero but grows as time evolves and peaks around horizon entry. We note that for cases without anisotropic stress (AD and CDI), $\zeta$ behaves qualitatively similarly to $\Phi$ or $\Psi$. However, $\zeta$ of DRDI is dramatically different than $\Phi$ or $\Psi$ due to the net anisotropic stress.
@@
where $k_{0}$ denotes the scale where the transition occurs. The $k^{3}$ dependence for small $k$ is generic for all modes that are outside the horizon around the generation of isocurvature (also known as white noise spectrum, see e.g., [Graham:2015rva, Irsic:2019iff, Feix:2020txt, Amin:2022nlh, Redi:2022llj, Elor:2023xbz, Graham:2024hah, deKruijf:2024voc, Ling:2024qfv, Amin:2025dtd]). If the observation data are mostly from this $k^{3}$ tail (i.e., for large $k_{0}$), the constraint can be applied to any model that has a $k^{3}$ spectrum up to a rescaling. The $k^{0}$ part for $k>k_{0}$ can originate from subhorizon dynamics during inflation, which naturally has nearly scale invariance. One concrete model that has both features is the non-thermal phase transition during inflation [Buckley:2024nen, Barir:2022kzo]. In this case, we can map the parameters in $\mathcal{P}(k)$ to model parameters. The amplitude $A_{\rm iso}\sim\Gamma_{\rm PT}/H_{\rm inf}^{4}$, where $\Gamma_{\rm PT}$ is the phase transition rate per volume and $H_{\rm inf}$ is the Hubble rate during inflation. The comoving scale when this phase transition starts is related to $k_{0}$.
We also show results for the log-normal power spectrum, which is widely studied for induced GWs [Pi:2020otn]. The log-normal spectrum is given as
@@
where $\Delta$ denotes the dimensionless width. This spectrum can be thought as a smeared version of the delta function spectrum, which can arise in certain inflationary models [Pi:2017gih, Kawasaki:1997ju, Frampton:2010sw, Cai:2018tuh].
@@
To source isocurvature perturbations, it generally needs new sources of perturbations other than the inflaton fluctuations. Additional scalar fields during inflation, such as axion or curvaton models (for reviews, see e.g., [Marsh:2015xka, Mazumdar:2010sa]), can source isocurvature after they decay. Other new dynamics like phase transitions during inflation can also excite isocurvature after they convert the vacuum energy of the phase transition into other species. If the decay or energy transfer is to the decoupled sector, such as CDM and DR, the isocurvature is straightforwardly inherited from the perturbations of the source. However, the BDI and NDI cases are subtler since baryons and neutrinos are tightly coupled to the SM bath and can efficiently transfer energy among them. The key point to maintain net isocurvature in baryons (neutrinos) is that the source must violate baryon (lepton) number, and there are no baryon or lepton number violation interactions in the thermal bath. Therefore the baryon (lepton) number is conserved and maintain the net isocurvature even they are in the thermal equilibrium. Such conditions can be achieved via a late decay of a particle that violates baryon or lepton number after the Electro-Weak sphaleron going out of the bath (when $T\sim 100$ GeV). For example, considering the thermal bath is initially adiabatic with no chemical potential in baryons (leptons), the number violating decays generate a net chemical potential $\mu$ in the phase space distribution of baryons (leptons) and opposite for antibaryons (antileptons). The total energy density of baryons and antibaryons differs from the case without chemical potential (adiabatic case) at $O(\mu^{2})$ for $\mu\ll T$, which becomes the isocurvature. Therefore, in these cases, the amplitude of isocurvature power spectrum is related to the second order of the chemical potential.
@@
  • As mentioned before, NDI is qualitatively different before and after neutrino decoupling. Remarkably, NDI is phenomenologically similar to CDI or BDI before decoupling, which reflects an interesting “conversion” between radiation and matter isocurvature.
@@ 3 Isocurvature Induced Gravitational Waves
With just scalar (metric and density) perturbations, tensor fluctuations $h_{ij}$ in Eq. (2.1) can not be sourced at the first order in perturbation theory due to the mismatch of the spin structure. However, at the second order, there is a source term for tensor modes from the combination of scalar perturbations. Following the standard formalism in Refs. [Kohri2018, Espinosa:2018eve, Pi:2020otn, Gong:2019mui], the equation of motion for tensor perturbation $h_{ij}$ (or gravitational waves) is given by
@@
where the source term is [Zhang:2022dgx]
$S_{ij}=$ $4{\cal P}^{ab}{}_{ij}\frac{1}{2}\left[-\partial_{a}\Phi\partial_{b}\Phi+\partial_{a}\Psi\partial_{b}\Phi+\partial_{a}\Phi\partial_{b}\Psi-3\partial_{a}\Psi\partial_{b}\Psi\right$.
$\left.-2\Phi\partial_{a}\partial_{b}\Phi-2\Psi\partial_{a}\partial_{b}\Psi+\frac{4}{3(1+w)}\partial_{a}(\Phi+{\cal H}^{-1}\Psi^{\prime})\partial_{b}(\Phi+{\cal H}^{-1}\Psi^{\prime})\right$.
$\left.-2\partial_{a}\partial_{b}(\Psi-\Phi)\Psi\right]~,$ (3.2)
where ${\cal P}^{ab}{}_{ij}$ is a transverse and traceless operator. For cases with non-vanishing anisotropic stress (e.g., DRDI), there will be additional contributions to the source term $S_{ij}$, which are proportional to the fraction of the density that has anisotropic stress [Weinberg:2003ur, Saga:2014jca, Zhang:2022dgx]. Since we only focus on $R_{\rm dr}\ll 1$ in this study, we neglect these contributions because they are further suppressed by $R_{\rm dr}$.
@@
with $\lambda$ denoting two polarizations of $h_{ij}$ and
$S_{\mathbf{k},\lambda}=4\int\frac{d^{3}\mathbf{q}}{(2\pi)^{3}}e^{ij}_{\mathbf{k},\lambda}q_{i}q_{j}{c(\mathbf{q})}c(|\mathbf{k}-\mathbf{q}|)f(\tau,q,|\mathbf{k}-\mathbf{q}|)~,$ (3.4)

### v2

@@ 1 Introduction
Gravitational waves (GWs) serve as another powerful probe of early-Universe physics. Because they free-stream, GWs can access information from much earlier times than the CMB photons. This prospect drives the continued, worldwide development of GW observatories designed to measure a broad range of frequencies. In particular, Pulsar Timing Arrays (PTA) provide a unique probe of GWs in the nano-Hertz frequency range. Several experiments such as NANOGrav [28], EPTA/InPTA [29], PPTA [30], CPTA [31] have recently observed a stochastic GW background and derived constraints for new physics. One possible source of GW that might contribute to PTA observations is called scalar induced GWs (SIGW) [^1] See Ref. [32] for a review. [33, 34, 35, 36, 37]: primordial scalar perturbations can source GWs at the second order in the perturbation theory. This framework has been applied to adiabatic perturbations and a certain type of isocurvature in cold dark matter (CDM) [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. In this work, we present a more comprehensive study of isocurvature induced GWs including four types of isocurvature: cold dark matter density isocurvature (CDI), baryon density isocurvature (BDI), neutrino density isocurvature (NDI) and free-streaming dark radiation density isocurvature (DRDI). We emphasize that, for the frequency range relevant for PTA observations, the modes entered the horizon before SM neutrino decoupling (when the temperature of the Universe is around MeV), and thus neutrinos are tightly coupled to the SM bath. Therefore, the initial conditions for NDI are qualitatively different from the standard ones where neutrinos are treated as free-streaming radiation. In this work, we derive novel initial conditions for all isocurvature modes with coupled neutrinos, and calculate the evolution of density and metric perturbations numerically. We find that the GW spectrum from NDI with coupled neutrinos is the same as that of CDI or BDI up to an overall coefficient, which leads to an interesting ‘conversion’ of isocurvature between radiation and matter sectors. Furthermore, DRDI results are dramatically different from those in the previously studied CDI case because of the presence of anisotropic stress. The size of DRDI induced GWs is significantly enhanced for high $k$ modes compared to that of CDI because the ratio of dark radiation (DR) density to the total bath is not suppressed at early times unlike the matter density. Therefore, PTA opens a new window to test BSM physics and places novel constraints via isocurvature induced GWs.
@@
To derive general constraints that can be applied to a broad class of models, we use two parameterizations of the primordial isocurvature power spectrum: a delta function and a broken power law [22]. Using the general formula for the source term of induced GWs, we obtain the spectrum of GWs and apply the data from NANOGrav collaboration to place constraints on the isocurvature power spectrum. Our results set stringent limits for wavenumbers around $10^{6}\,\text{Mpc}^{-1}$, which are complementary to other cosmological constraints on large scales.
@@ 2 Isocurvature Initial Conditions Before Neutrino Decoupling
To study perturbations in cosmology [54, 55, 56], we start with the perturbed FRW metric in the conformal Newtonian gauge,
@@
Now we move on to study different types of perturbation in cosmology. The gauge invariant curvature perturbation can be written in the Newtonian gauge as [3]
@@
where $R_{i}\equiv\bar{\rho}_{i}/\bar{\rho}_{r}$ with $\bar{\rho}_{r}$ being the total background radiation density and $i$ denotes species in the radiation bath. The gauge-invariant isocurvature perturbation, or the relative entropy perturbation is defined as
${\cal S}_{\alpha\beta}=3(\zeta_{\alpha}-\zeta_{\beta})=\frac{\delta_{\alpha}}{1+w_{\alpha}}-\frac{\delta_{\beta}}{1+w_{\beta}}~,$ (2.8)
where $w\equiv\bar{p}/\bar{\rho}$ with $\bar{p}$ being the background momentum density.
@@
All other solutions are called isocurvature initial conditions, which have vanishing curvature perturbation $\zeta=0$ but at least one non-zero $\cal S_{\alpha\beta}$. Since there are multiple species in the Universe, we can define different isocurvature initial conditions by different choices of non-vanishing $\cal S_{\alpha\beta}$. In this paper, we choose the adiabatic SM radiation bath $\gamma_{*}$ as the reference and we consider only one non-vanishing $\mathcal{S}_{\alpha\gamma_{*}}$ in each case as $\tau\to 0$. [^2] Since $\mathcal{S}_{\alpha\gamma_{*}}$ is always the same as $\mathcal{S}_{\alpha\gamma}$, it is equivalent to choose the photon bath as the reference. For example, NDI is defined by setting $\mathcal{S}_{\alpha\gamma_{*}}=0$ except for $\mathcal{S}_{\nu\gamma_{*}}\neq 0$. Similarly, we can define BDI ($\mathcal{S}_{b\gamma_{*}}\neq 0$), CDI ($\mathcal{S}_{c\gamma_{*}}\neq 0$), and DRDI ($\mathcal{S}_{\textrm{dr}\gamma_{*}}\neq 0$). Once fixing the value of the non-zero $\mathcal{S}_{\alpha\gamma_{*}}$ (e.g., we choose $\mathcal{S}_{c\gamma_{*}}=1$ for CDI), initial conditions for all metric and density perturbations are uniquely determined. We note that the definition of isocurvature initial conditions may vary depending on the choice of non-vanishing $\cal S_{\alpha\beta}$. Another commonly studied choice is based on $\mathcal{S}_{\alpha r}$ which involves the perturbation of the radiation bath [2]. We find this choice matches our definition if the isocurvature component is not in the radiation bath like BDI and CDI, but differs for NDI and DRDI. Since $\zeta_{r}$ in the NDI case contains neutrino perturbations that have isocurvature, setting $\mathcal{S}_{br/cr}=0$ does not respect the adiabaticity of baryons and CDM. Therefore, for the study of new physics that generate isocurvature in certain species and the rest remain adiabatic, it is more convenient to use the definition based on $\mathcal{S}_{\alpha\gamma_{*}}$ instead of $\mathcal{S}_{\alpha r}$.
@@
As mentioned before, we separate $\nu$ from $\gamma_{*}$ for NDI. Setting $\mathcal{S}_{\nu\gamma_{*}}=3/4$, [^3] This is the convention commonly used for radiation isocurvature, which ensures the leading isocurvature perturbation (e.g., $\delta_{\nu}$ in NDI) is normalized close to unity [2]. we get perturbations in NDI as
@@
$\theta_{c}$ $=$ $\frac{1}{3}\theta_{b}=\frac{1}{3}\theta_{\gamma_{*}}=\frac{1}{3}\theta_{\nu}=-\frac{1}{24}R_{\nu}\omega k^{2}\tau^{2}.$ (2.13)
Although $\delta_{\alpha}$ approaches a constant as $\tau\to 0$, they do not source metric or curvature perturbations because the total density perturbation $\delta=(\sum_{\alpha}\bar{\rho}_{\alpha}\delta_{\alpha})/\bar{\rho}$ vanishes as $\tau\to 0$.
@@
$\theta_{c}$ $=$ $-\frac{(1-R_{\rm dr})R_{\rm dr}}{15+4R_{\rm dr}}k^{2}\tau$
@@
We show the time evolution of $\Psi$ and $\Phi$ for different cases in the left panel of figure 1. For illustrative purposes, we just show the AD, CDI and DRDI cases. Since initial conditions for CDI is qualitatively similar to BDI and NDI up to an overall rescaling as mentioned before, these choices covers all relevant features. As seen from figure 1, $\Psi=\Phi\propto C$ for AD and $\Psi=\Phi\propto\omega\tau$ for CDI at superhorizon. They both become damped oscillations after entering the horizon. Around horizon crossing ($k\tau\sim 1$), the CDI case is suppressed by $\omega/k$ (or $k_{\rm eq}/k$ because $k_{\rm eq}=\sqrt{2}\omega$) for large $k$. Therefore we rescale the CDI case by the inverse of this factor. The DRDI case, however, has qualitatively different $\Phi$ and $\Psi$ evolutions due to the presence of anisotropic stress from free-streaming DR. In DRDI, $\Phi,\Psi,\Phi-\Psi\propto R_{\rm dr}$ (see Eq. (2.16)) at superhorizon. They both start damped oscillations after horizon entry and become the same deep inside the horizon due to the anisotropic stress being suppressed (see Eq. (2.15)). Moreover, we present the evolution of the curvature perturbation $\zeta$ in the right panel of figure 1. $\zeta$ of AD starts at one by definition and oscillates after horizon entry. For isocurvature cases, $\zeta$ starts at zero but grows as time evolves and oscillates after horizon entry. We note that for cases without anisotropic stress (AD and CDI), $\zeta$ behaves qualitatively similarly to $\Phi$ or $\Psi$. However, $\zeta$ of DRDI is dramatically different than $\Phi$ or $\Psi$ due to the net anisotropic stress.
@@
where $k_{0}$ denotes the scale where the transition occurs. The $k^{3}$ dependence for small $k$ is generic for all modes that are outside the horizon around the generation of isocurvature (also known as white noise spectrum, see e.g., [57, 58, 59, 60, 61, 62, 63, 64, 65, 66]). If the observation data are mostly from this $k^{3}$ tail (i.e., for large $k_{0}$), the constraint can be applied to any model that has a $k^{3}$ spectrum up to a rescaling. The $k^{0}$ part for $k>k_{0}$ can originate from subhorizon dynamics during inflation, which naturally has nearly scale invariance. One concrete model that has both features is the non-thermal phase transition during inflation [67, 68]. In this case, we can map the parameters in $\mathcal{P}(k)$ to model parameters. The amplitude $A_{\rm iso}\sim\Gamma_{\rm PT}/H_{\rm inf}^{4}$, where $\Gamma_{\rm PT}$ is the phase transition rate per volume and $H_{\rm inf}$ is the Hubble rate during inflation. The comoving scale when this phase transition starts is related to $k_{0}$.
We also show results for the log-normal power spectrum, which is widely studied for induced GWs [69]. The log-normal spectrum is given as
@@
where $\Delta$ denotes the dimensionless width. This spectrum can be thought as a smeared version of the delta function spectrum, which can arise in certain inflationary models [70, 71, 72, 73].
@@
To source isocurvature perturbations, it generally needs new sources of perturbations other than the inflaton fluctuations. Additional scalar fields during inflation, such as axion or curvaton models (for reviews, see e.g., [74, 75]), can source isocurvature after they decay. Other new dynamics like phase transitions during inflation can also excite isocurvature after they convert the vacuum energy of the phase transition into other species. If the decay or energy transfer is to the decoupled sector, such as CDM and DR, the isocurvature is straightforwardly inherited from the perturbations of the source. However, the BDI and NDI cases are subtler since baryons and neutrinos are tightly coupled to the SM bath and can efficiently transfer energy among them. The key point to maintain net isocurvature in baryons (neutrinos) is that the source must violate baryon (lepton) number, and there are no baryon or lepton number violation interactions in the thermal bath. Therefore the baryon (lepton) number is conserved and maintain the net isocurvature even they are in the thermal equilibrium. Such conditions can be achieved via a late decay of a particle that violates baryon or lepton number after the Electro-Weak sphaleron going out of the bath (when $T\sim 100$ GeV). For example, considering the thermal bath is initially adiabatic with no chemical potential in baryons (leptons), the number violating decays generate a net chemical potential $\mu$ in the phase space distribution of baryons (leptons) and opposite for antibaryons (antileptons). The total energy density of baryons and antibaryons differs from the case without chemical potential (adiabatic case) at $O(\mu^{2})$ for $\mu\ll T$, which becomes the isocurvature. Therefore, in these cases, the amplitude of isocurvature power spectrum is related to the second order of the chemical potential.
@@
  • As mentioned before, NDI is qualitatively different before and after neutrino decoupling. Remarkably, NDI is phenomenologically similar to CDI or BDI before decoupling, which reflects an interesting “conversion” between radiation and matter isocurvature. Transition between NDI before and after neutrino decoupling is non-trivial, which requires to study the full Boltzmann equations with proper time-dependent collision terms (see e.g., Refs. [76, 77]). Detailed study of this transition is beyond the scope of current work.
@@ 3 Isocurvature Induced Gravitational Waves
With just scalar (metric and density) perturbations, tensor fluctuations $h_{ij}$ in Eq. (2.1) can not be sourced at the first order in perturbation theory due to the mismatch of the spin structure. However, at the second order, there is a source term for tensor modes from the combination of scalar perturbations. The equation of motion for tensor fluctuation $h_{ij}$ can be derived from the second-order Einstein’s equation, which depends on the second-order energy momentum tensor $T^{(2)}_{ij}$. For cases without anisotropic stress (AD/CDI/BDI/NDI), the whole radiation bath is strongly coupled. Under the tight-coupling approximation, the leading contribution to the second-order energy momentum tensor is given as [78, 79]
$T^{(2)}_{ij}=\frac{4}{3}\bar{\rho}_{r}\partial_{i}v_{\gamma_{*}}\partial_{j}v_{\gamma_{*}}+(\textrm{diagonal part})\delta_{ij}\,~~~(\textrm{without DR}),$ (3.1)
where $v$ is another way to denote the velocity perturbation, which is related to the previously defined $\theta$ as $\theta(k)=-k^{2}v(k)$. [^4] We drop the superscript label (1) for first order perturbations for simplicity. Here we neglect contributions from matter components (CDM and baryons) because they are highly suppressed deep inside the radiation domination. For the DRDI case, free-streaming DR cannot be treated as a coupled fluid, and thus the second-order energy momentum tensor reads
$T^{(2)}_{ij}=\frac{4}{3}\bar{\rho}_{\gamma_{*}}\partial_{i}v_{\gamma_{*}}\partial_{j}v_{\gamma_{*}}+\bar{\rho}_{\textrm{dr}}\,\Pi^{(2)}_{\textrm{dr}\,ij}+(\textrm{diagonal part})\delta_{ij}\,~~~(\textrm{DRDI}),$ (3.2)
where $\Pi^{(2)}_{\textrm{dr}\,ij}$ denotes the contribution of DR to the second-order energy momentum tensor. $\Pi^{(2)}_{\textrm{dr}\,ij}$ is determined by the Boltzmann equation at the second order, which is not simply $\propto\partial_{i}v_{\textrm{dr}}\partial_{j}v_{\textrm{dr}}$ (see Refs. [78, 79, 80, 81] and appendix C).
Following the standard formalism in Refs. [82, 83, 69, 84], the equation of motion for tensor perturbation $h_{ij}$ (or gravitational waves) is given by
@@
where the source term for AD/CDI/BDI/NDI cases in the radiation dominated Universe is [32]
$S_{ij}=$ $4{\cal P}^{ab}{}_{ij}\left[\partial_{a}\Phi\partial_{b}\Phi+4\pi Ga^{2}\frac{4}{3}\bar{\rho}_{r}\partial_{a}v_{\gamma_{*}}\partial_{b}v_{\gamma_{*}}\right]~~~~(\textrm{without DR}),$ (3.4)
where ${\cal P}^{ab}{}_{ij}$ is a transverse and traceless operator and we used $\Psi=\Phi$ due to the absence of anisotropic stress. We note that NDI has the same expression because $\bar{\rho}_{r}=\bar{\rho}_{\gamma_{*}}+\bar{\rho}_{\nu}$ and $v_{\gamma_{*}}=v_{\nu}$.
@@
with $\lambda=+,\times$ denoting two polarizations of $h_{ij}$ and
$S_{\mathbf{k},\lambda}=4\int\frac{d^{3}\mathbf{q}}{(2\pi)^{3}}e^{ij}_{\mathbf{k},\lambda}q_{i}q_{j}{c(\mathbf{q})}c(\mathbf{k}-\mathbf{q})f(\tau,q,|\mathbf{k}-\mathbf{q}|)~,$ (3.6)